# Prediction of Protein Three-dimensional Structures via a Hardware-Executable Quantum Computing Framework

### v1

@@ Abstract
Accurate prediction of protein active site structures remains a central challenge in structural biology, particularly for short and flexible peptide fragments where conventional methods often fail. Here, we present a quantum computing framework specifically developed for utility-level quantum processors to address this problem. Starting from an amino acid sequence, we formulate the structure prediction task as a ground-state energy minimization problem using the Variational Quantum Eigensolver (VQE). Amino acid connectivity is encoded on a tetrahedral lattice model, and structural constraints—including steric, geometric, and chirality terms—are mapped into a problem-specific Hamiltonian expressed as sparse Pauli operators. The optimization is executed via a two-stage architecture separating energy estimation and measurement decoding, allowing noise mitigation under realistic quantum device conditions. We evaluate the framework on 23 randomly selected real protein fragments from the PDBbind dataset, as well as 7 real fragments from proteins with therapeutic potential, and run the experiments on the IBM–Cleveland Clinic quantum processor. Structural predictions are benchmarked against AlphaFold3 (AF3) using identical postprocessing and docking procedures. Our quantum method outperformed AF3 in both RMSD (Root-Mean-Square Deviation) and docking efficacy. This work demonstrates, for the first time, a complete end-to-end pipeline for biologically relevant structure prediction on real quantum hardware, highlighting its engineering feasibility and practical advantage over existing classical and deep learning approaches.
@@ Introduction
Predicting the local three-dimensional structures of proteins, particularly at ligand-binding sites, is a central problem in structural biology with broad implications for molecular function annotation, drug design, and protein engineering[1]. The conformational landscape of these regions is often characterized by high flexibility and functional specificity, making accurately model structures both scientifically critical and technically challenging[2]. Traditional physics-based methods, such as molecular dynamics (MD), can in principle sample native-like structures but suffer from prohibitive computational costs and scalability issues[3, 4]. On the other hand, recent breakthroughs in deep learning, such as AlphaFold2 and AlphaFold3, have demonstrated remarkable success in full-length protein structure prediction[5, 6]. However, their performance substantially deteriorates when applied to short peptide fragments or highly flexible local domains, where limited sequence context and sparse evolutionary information lead to structural ambiguities and reduced accuracy [7]. In particular, short protein fragments involved in ligand binding often contain fewer than 20 amino acids and exhibit complex energetic constraints that are difficult to capture with purely data-driven models. Most of the results predicted by deep learning methods are excellent; however, in some cases, issues such as overfitting or incorrect predictions may occur. Deep learning approaches still have certain blind spots, particularly in the prediction of short fragments [8]. Although several methods (e.g. modeling the problem as a binary optimization task for optimization, employing quantum annealing algorithms, or using the QAOA (Quantum Approximate Optimization Algorithm) method) have previously been proposed for protein structure prediction using quantum computing, their validation in real-world scenarios remains highly limited. Some approaches are only capable of generating predictions, but the performance of these predictions often falls short of being practically applicable to real problems (e.g. the docking application discussed in this paper) [9]. Therefore, there is an increasing need for protein prediction frameworks that operate under limited input information capacity (e.g., short protein fragments), make efficient use of quantum resources, and deliver high-precision results to meet the requirements of tasks such as docking. Quantum computing offers a new paradigm for addressing classically intractable optimization problems by exploiting quantum parallelism, entanglement, and interference. While current quantum hardware remains limited by noise, decoherence, and qubit count—commonly referred to as utility-level processors—it has reached a level of maturity that enables real-world applications in small-scale, high-precision tasks. Among hybrid quantum-classical algorithms, the Variational Quantum Eigensolver (VQE) is particularly well-suited to molecular energy landscape optimization [10]. By encoding molecular constraints into parameterized quantum circuits, VQE enables the estimation of ground-state energies for quantum Hamiltonians constructed from physical models.
@@
In this work, we present an end-to-end quantum computing framework for predicting the local structure of protein binding site fragments using utility-level quantum hardware. Our approach begins with an amino acid sequence and maps its spatial configuration to a tetrahedral lattice, where each residue is represented as a node with four directional connectivity options. These configurations are encoded as quantum measurement states corresponding to discrete Hamiltonians that incorporate steric constraints, chirality terms, geometric restrictions, and amino acid interactions. To mitigate the effects of hardware noise, we design a two-stage execution architecture that separates energy estimation from structural decoding. This strategy enhances execution stability while maintaining fidelity in structural recovery. We validate our framework using 30 short peptide sequences extracted from the PDBbind database [11, 12, 13]. The sequences are processed through our quantum pipeline and executed on the IBM-Cleveland Clinic 127-qubit superconducting quantum processor. For benchmarking, we compare the resulting structures against those predicted by AlphaFold3, followed by uniform postprocessing and ligand docking using AutoDock[14, 15, 16, 17, 6]. Our quantum method outperforms AlphaFold3 in both docking success rate and RMSD (Root-Mean-Square Deviation), demonstrating the feasibility and effectiveness of quantum structure prediction on real quantum hardware. This study marks one of the first demonstrations of quantum computing applied to biologically relevant structure prediction on real devices. It establishes a blueprint for leveraging utility-level quantum processors in solving domain-specific, engineering-scale problems that remain out of reach for both classical simulations and data-driven models.
@@ A Quantum Engineering Framework for Structure Prediction
To validate the executability and engineering robustness of the proposed framework on real quantum hardware, we deployed the full structure prediction pipeline on the IBM-Cleveland Clinic’s 127-qubit superconducting quantum processor. All tasks were executed directly on physical hardware rather than in simulation environments. Depending on the fragment length and the complexity of encoded amino acid interactions, the number of active qubits required for each task ranged from several dozen to over a hundred. We used a tetrahedral lattice model to discretize the amino acid sequence and constructed a sparse Pauli Hamiltonian incorporating steric repulsion, chirality constraints, geometric feasibility, and residue-level interaction energies. During the variational optimization phase, we employed a parameterized quantum circuit composed of native single-qubit rotation gates (e.g., $\text{Ry}(\theta)$) and controlled-NOT (CX) gates. These circuits were compiled and topology-mapped using Qiskit’s transpiler to ensure compatibility with the IBM-Cleveland Clinic quantum processor. To improve tolerance to hardware noise and maintain structural prediction stability, we adopted a two-stage execution strategy, architecturally separated into a variational optimization module and a dedicated measurement module. In the first stage, the Variational Quantum Eigensolver (VQE) was used to minimize the expectation value of the constructed Hamiltonian. Each energy evaluation was performed using 2,000 shots for quantum energy estimation, combined with at least 200 iterations of the COBYLA (Constrained Optimization BY Linear Approximations) classical optimizer. In the second stage, the optimized parameter set from the first phase was fixed and used to reconstruct a non-parametric circuit, which was then recompiled into a stable quantum circuit composed entirely of native gates. This fixed circuit was executed 20,000 times to obtain measurement bitstrings representing structural states. This two-phase process ensured reproducibility and robustness under noise-constrained physical environments. For testing, we selected 23 representative peptide fragments from the PDBbind database [11, 12, 13]. Each fragment originated from a distinct protein-ligand complex with experimentally determined receptor structures. These fragments were chosen to be structurally independent and diverse in both residue composition and sequence ordering, enabling us to assess the generalization capability of our framework across varying input conditions. Table 1 summarizes the selected samples’ PDB IDs, residue sequences, sequence lengths, and execution parameters including number of qubits, circuit depth, execution time, and the system’s minimum, maximum, and energy range during optimization. Due to the high computational cost of quantum hardware, we adopted a strategy of low-iteration VQE optimization coupled with high-precision measurements. If resource constraints were relaxed, it would be feasible to achieve better convergence paths and structure prediction accuracy by increasing the number of optimization steps and adopting gradient-based optimizers. The reason why the execution times shown in the table do not increase proportionally with the length of the protein fragments is mainly due to the current task scheduling mechanism on quantum computers. Since the scheduling strategy is not yet fully intelligent, some tasks may be preempted by higher-priority jobs—such as those submitted by accounts with higher access levels or with shorter estimated runtimes—which results in additional session delays and idle time.
@@
The 23 test cases ranged from short peptides with 5 amino acids to medium-length chains of 10–14 residues, ensuring the framework’s compatibility across varying sequence lengths. Several clear trends were observed from the collected execution data: first, the number of qubits increased significantly with sequence length, as more ancillary qubits are required to represent growing residue-residue interactions and constraints. Second, the system energy exhibited a stepwise increase as sequence length increased, with sharp jumps observed at certain threshold lengths. Using the system’s energy range as an indicator, we found that increasing the sequence length from 6 to 7 residues led to a jump in energy range from approximately 260 a.u. (atomic units, Hartree energy) to over 600 a.u., an increase of about 131%. From 7 to 9 residues, energy range increased from around 610 a.u. to approximately 1,350 a.u. (+121%), and further to over 2,100 a.u. at 10 residues (+56%). Between 10 and 11 residues, energy range continued to rise to 2,500–2,700 a.u. (+20–30%), then reached 4,600–5,300 a.u. at 13 residues (+85%), and finally surpassed 9,000 a.u. at 14 residues (+72% over 13 residues and +2,700% over the initial 6-residue case). Both the minimum and maximum energy values also expanded rapidly with sequence length. For example, minimum energy increased from approximately 10 a.u. (5 residues) to over 23,000 a.u. (14 residues), while maximum energy increased from  15 a.u. to over 32,000 a.u.—representing increases of approximately 230,000% and 210,000%, respectively. These nonlinear surges demonstrate the exponential growth of the quantum energy landscape as conformational dimensionality increases. This nonlinear energy amplification reveals two key insights. First, the quantum system is capable of encoding subtle structural variations with high precision: even the addition or removal of a single amino acid causes energy changes distinguishable by the quantum model. Second, the energy of the quantum system, as encoded under our Hamiltonian mapping, is highly correlated with the underlying conformational energy of the protein fragment. Thus, the lowest-energy quantum state produced by VQE directly corresponds to the predicted structure, validating the soundness of our prediction strategy. It is precisely due to this energy amplification behavior that we modularized the measurement stage, allowing it to operate independently with non-parametric circuits for high-precision bitstring sampling. This design preserves prediction fidelity while maintaining engineering tractability under realistic hardware constraints.
@@
4aoi VVLPYMKHGDLRNF 14 1155–1168 102 23245.37 32378.95 9133.58 13328.65 413
4cig VRDQAEHLKTAVQM 14 165–178 102 21375.59 29846.54 8470.94 17293.54 413
4tmk IEGLEGAGKTTARN 14 8–21 102 22590.21 29135.42 6545.21 199292.66 413
3d7z YLVTHLMGADLNNI 14 103–116 102 22979.86 29707.30 6727.43 156289.48 413
4jpx DYLEAYGKGGVKA 13 154–166 92 16962.09 22231.95 5269.86 90422.62 373
1yc4 ELISNSSDALDKI 13 47–59 92 16129.38 20745.81 4616.42 15777.29 373
5cqu RKLGRGKYSEVFE 13 43–55 92 17865.39 22801.51 4936.12 7620.94 373
2qbs HCSAGIGRSGT 11 214–224 72 6691.57 9356.87 2665.30 13899.11 293
4f5y GLAWSYYIGYL 11 158–168 72 6408.50 8858.60 2450.10 6157.46 293
5cxa FDGKGGILAHA 11 174–184 72 6946.43 9298.82 2352.40 5638.71 293
4mc1 LLDTGADDTV 10 23–32 63 4092.24 6199.23 2106.99 5609.02 257
3d83 YLVTHLMGAD 10 103–112 63 4235.34 6119.16 1883.82 19833.57 257
3vf7 LLDTGADDTV 10 23–32 63 3975.02 6162.42 2187.40 5348.25 257
4y79 DACQGDSGG 9 189–197 54 1549.16 2874.21 1325.05 207445.70 221
1e2l AQITMGMPY 9 124–132 54 1509.66 2837.82 1328.15 12951.69 221
1gx8 SAPLRVYVE 9 36–44 54 1626.02 3053.53 1427.51 14080.77 221
1m7y TAGATSANE 9 117–125 54 1420.38 2714.98 1294.60 12918.04 221
3dx3 HNDPGWI 7 90–96 38 339.99 962.62 622.63 4661.24 157
5c28 CDLCSVT 7 663–669 38 386.81 792.78 405.97 114029.96 157
3ibi IQFHFH 6 91–96 23 120.66 455.42 334.76 4486.62 97
2v25 ATFTIT 6 81–86 23 100.42 340.83 240.42 22356.46 97
3ckz VKDRS 5 149–153 12 10.43 14.65 4.22 5763.36 53
3eax RYRDV 5 45–49 12 10.36 16.02 5.66 4028.72 53
@@
Our RMSD-based geometric comparison across 23 protein binding site fragments indicates that the quantum method achieved lower RMSD values in 18 samples, outperforming AlphaFold3 (AF3), which had better results in the remaining 5 cases (Figure S1). The average RMSD for quantum predictions was 3.33 Å, with a median of 3.53 Å, while AF3 predictions yielded an average RMSD of 3.87 Å and a median of 3.92 Å. Notably, the quantum method exhibited stronger structural fidelity in shorter sequences (5-9 residues) and in regions with higher structural flexibility. The predicted structures were used as ligands in docking experiments with a standardized receptor structure, and binding affinity were recorded over multiple runs. Results demonstrate that quantum-predicted structures achieved better binding affinities in 21 out of 23 cases (Figure S2). The average binding affinity for the quantum method was -4.38 kcal/mol, while the AF3-predicted structures averaged -4.00 kcal/mol. This suggests that quantum predictions not only exhibit superior geometric precision, but also more accurately capture the energetic characteristics of real biomolecular interactions. Furthermore, a cross-analysis of structural and functional metrics reveals that in most cases, quantum-predicted structures simultaneously exhibit lower RMSD and more favorable binding affinities, indicating a strong correlation between geometric and energetic consistency. This pattern suggests that, within the current problem scale, quantum predictions effectively capture both low-energy conformations and biologically relevant binding configurations. In summary, the quantum computing framework presented in this study delivers structural outputs on real quantum hardware that outperform state-of-the-art deep learning methods in both geometric and functional aspects. Particularly in low-information or small-sample prediction scenarios, where traditional models tend to degrade, our framework demonstrates significant advantages, highlighting its potential utility and scalability in structural biology and bioengineering applications.
@@
Compared to AlphaFold3 predictions, the quantum method’s predicted structures achieved approximately 20% reductions in the upper and lower bounds of RMSD in different test cases during docking tests, indicating higher consistency and accuracy in structural prediction (Figure 3H-I). In terms of affinity score, the quantum method achieved an average improvement of about 10% over AlphaFold3 (Figure 3G). Additionally, the quantum method exhibited a tighter distribution of affinity scores, with significantly fewer outliers, further validating its reliability in optimizing docking performance. For protein structure prediction, besides evaluating predictive accuracy in docking tasks through binding affinity scores and RMSD bounds of docking poses, another important metric is the RMSD itself. The RMSD (root-mean-square deviation) metric is used to quantify structural differences between the predicted and experimental structures. In our experiments, we compared the RMSD of structures predicted by the VQE method and those predicted by AF3 against the corresponding experimental structures in the dataset. As shown in Figure 3F, the VQE method outperformed AF3 in six out of seven cases. For 6mu3, due to its short chain length and the limited degrees of freedom of the modeling approach, it lagged behind AF3 by only a percentage-level margin. Although most experimental results did not significantly surpass AF3 predictions due to limitations in modeling precision, they nevertheless demonstrate that quantum methods possess certain advantages.

### v2

# A Quantum Framework for Protein Binding-Site Structure Prediction on Utility-Level Quantum Processors

@@ Abstract
Accurate prediction of protein active site structures remains a central challenge in structural biology, particularly for short and flexible peptide fragments where conventional and simulation-based methods often fail. Here, we present a quantum computing framework specifically developed for utility-level quantum processors to address this problem. Starting from an amino acid sequence, we formulate the structure prediction task as a ground-state energy minimization problem using the Variational Quantum Eigensolver (VQE). Amino acid connectivity is encoded on a tetrahedral lattice model, and structural constraints—including steric, geometric, and chirality terms—are mapped into a problem-specific Hamiltonian expressed as sparse Pauli operators. The optimization is executed via a two-stage architecture separating energy estimation and measurement decoding, allowing noise mitigation under realistic quantum device conditions. We evaluate the framework on 23 randomly selected real protein fragments from the PDBbind dataset, as well as 7 fragments from therapeutically relevant proteins, and run the experiments on the IBM–Cleveland Clinic quantum processor. Structural predictions are benchmarked against both AlphaFold3 (AF3) and classical simulation–based approaches using identical postprocessing and docking procedures. Our quantum method outperformed both AF3 and classical models in RMSD (Root-Mean-Square Deviation) and docking efficacy. This work demonstrates, for the first time, a complete end-to-end pipeline for biologically relevant structure prediction on real quantum hardware, highlighting its engineering feasibility and practical advantage over existing classical and deep learning approaches.
@@ Introduction
Predicting the local three-dimensional structures of proteins, particularly at ligand-binding sites, is a central challenge in structural biology with broad implications for molecular function annotation, drug design, and protein engineering [42]. The conformational landscape of these regions is often characterized by high flexibility and functional specificity, making accurate modeling both scientifically critical and technically challenging [54]. Traditional physics-based methods, such as molecular dynamics (MD), can in principle sample native-like structures but suffer from prohibitive computational costs and scalability issues [27]. By contrast, recent breakthroughs in deep learning, exemplified by AlphaFold2 and AlphaFold3, have achieved remarkable success in full-length protein structure prediction [32, 1]. However, their performance substantially deteriorates when applied to short peptide fragments or highly flexible local domains, where limited sequence context and sparse evolutionary information lead to structural ambiguities and reduced accuracy [64, 5]. In particular, short fragments that constitute ligand-binding pockets typically contain fewer than 20 amino acids and exhibit complex energetic constraints that are difficult to capture with purely data-driven models.
While several approaches have explored quantum computing for protein structure prediction—including formulations as binary optimization tasks, quantum annealing schemes, and the Quantum Approximate Optimization Algorithm (QAOA) [59]—their validation in realistic, biologically relevant settings remains limited. Most prior methods have been evaluated only on toy examples or simulations and often lack the accuracy needed for downstream applications such as ligand docking. This gap underscores the need for frameworks that can operate under limited input information (e.g., short fragments), make efficient use of quantum resources, and deliver predictions of sufficient precision to support molecular docking and design. Quantum computing offers a fundamentally new paradigm for tackling such problems by exploiting quantum parallelism, entanglement, and interference. Although current hardware is constrained by noise, decoherence, and limited qubit counts—commonly referred to as utility-level processors—it has reached a level of maturity that enables early-stage, domain-specific applications. Among hybrid quantum-classical algorithms, the Variational Quantum Eigensolver (VQE) is particularly well-suited for energy landscape optimization [51, 76]. By encoding physical constraints into parameterized quantum circuits, VQE enables estimation of ground-state energies for Hamiltonians constructed from simplified structural models.
@@
In this work, we present an end-to-end quantum computing framework for predicting the local structure of protein binding-site fragments on utility-level quantum hardware. Our approach begins with an amino acid sequence and maps its spatial configuration to a tetrahedral lattice, where each residue is represented as a node with four directional connectivity options. These configurations are encoded as quantum measurement states corresponding to discrete Hamiltonians that incorporate steric constraints, chirality terms, geometric restrictions, and amino acid interactions. To mitigate the effects of hardware noise, we design a two-stage execution architecture that separates energy estimation from structural decoding. This strategy enhances execution stability while preserving fidelity in structural recovery. We validate our framework using short peptide sequences extracted from the PDBbind database [73, 74, 44] and execute them on the IBM–Cleveland Clinic 127-qubit superconducting quantum processor. For benchmarking, we compare our results against both AlphaFold3 and classical simulation–based prediction methods, followed by uniform postprocessing and docking evaluation using AutoDock [30, 20, 71, 16]. Across all test fragments, the quantum framework consistently achieves lower RMSD values and more favorable docking affinities than both AlphaFold3 and classical models, thereby demonstrating the feasibility and practical advantage of quantum-assisted structural prediction on real devices.
The key contributions of this work are as follows: (1) We introduce a coarse-grained, vector-based lattice encoding and corresponding Hamiltonian formulation tailored for short peptide fragments at binding pockets, which reduces side-chain complexity and produces a smoother, noise-resilient energy landscape. (2) We develop an end-to-end quantum computing framework that separates energy estimation from structural decoding, thereby enhancing robustness against hardware noise. (3) We demonstrate, for the first time, biologically relevant structure prediction for docking-oriented fragments on a real 127-qubit superconducting quantum processor, achieving superior RMSD and docking performance compared to both AlphaFold3 and classical simulation–based approaches. Together, these advances highlight a practical role for today’s noisy quantum hardware in structural biology and establish a blueprint for integrating quantum computing into protein docking and design workflows.
@@ A Quantum Engineering Framework for Structure Prediction
To validate the executability and engineering robustness of the proposed framework on real quantum hardware, we deployed the full structure prediction pipeline on the IBM-Cleveland Clinic’s 127-qubit superconducting quantum processor. All tasks were executed directly on physical hardware rather than in simulation environments. Depending on the fragment length and the complexity of encoded amino acid interactions, the number of active qubits required for each task ranged from several dozen to over a hundred. We used a tetrahedral lattice model to discretize the amino acid sequence and constructed a sparse Pauli Hamiltonian incorporating steric repulsion, chirality constraints, geometric feasibility, and residue-level interaction energies. During the variational optimization phase, we employed a parameterized quantum circuit composed of native single-qubit rotation gates (e.g., $\text{Ry}(\theta)$) and controlled-NOT (CX) gates. These circuits were compiled and topology-mapped using Qiskit’s transpiler to ensure compatibility with the IBM-Cleveland Clinic quantum processor. To improve tolerance to hardware noise and maintain structural prediction stability, we adopted a two-stage execution strategy, architecturally separated into a variational optimization module and a dedicated measurement module. In the first stage, the Variational Quantum Eigensolver (VQE) was used to minimize the expectation value of the constructed Hamiltonian. Each energy evaluation was performed using 2,000 shots for quantum energy estimation, combined with at least 200 iterations of the COBYLA (Constrained Optimization BY Linear Approximations) classical optimizer. In the second stage, the optimized parameter set from the first phase was fixed and used to reconstruct a non-parametric circuit, which was then recompiled into a stable quantum circuit composed entirely of native gates. This fixed circuit was executed 20,000 times to obtain measurement bitstrings representing structural states. This two-phase process ensured reproducibility and robustness under noise-constrained physical environments. For testing, we selected 23 representative peptide fragments from the PDBbind database [44]. Each fragment originated from a distinct protein-ligand complex with experimentally determined receptor structures. These fragments were chosen to be structurally independent and diverse in both residue composition and sequence ordering, enabling us to assess the generalization capability of our framework across varying input conditions. Table 1 summarizes the selected samples’ PDB IDs, residue sequences, sequence lengths, and execution parameters including number of qubits, circuit depth, execution time, and the system’s minimum, maximum, and energy range during optimization. Due to the high computational cost of quantum hardware, we adopted a strategy of low-iteration VQE optimization coupled with high-precision measurements. If resource constraints were relaxed, it would be feasible to achieve better convergence paths and structure prediction accuracy by increasing the number of optimization steps and adopting gradient-based optimizers. The reason why the execution times shown in the table do not increase proportionally with the length of the protein fragments is mainly due to the current task scheduling mechanism on quantum computers. Since the scheduling strategy is not yet fully intelligent, some tasks may be preempted by higher-priority jobs—such as those submitted by accounts with higher access levels or with shorter estimated runtimes—which results in additional session delays and idle time.
@@
The 23 test cases ranged from short peptides with 5 amino acids to medium-length chains of 10–14 residues, ensuring the framework’s compatibility across varying sequence lengths. Several clear trends were observed from the collected execution data: first, the number of qubits increased significantly with sequence length, as more ancillary qubits are required to represent growing residue-residue interactions and constraints. Second, the system energy exhibited a stepwise increase as sequence length increased, with sharp jumps observed at certain threshold lengths. Using the system’s energy range as an indicator, we found that increasing the sequence length from 6 to 7 residues led to a jump in energy range from approximately 260 a.u. (atomic units, Hartree energy) to over 600 a.u., an increase of about 131%. From 7 to 9 residues, energy range increased from around 610 a.u. to approximately 1,350 a.u. (+121%), and further to over 2,100 a.u. at 10 residues (+56%). Between 10 and 11 residues, energy range continued to rise to 2,500–2,700 a.u. (+20–30%), then reached 4,600–5,300 a.u. at 13 residues (+85%), and finally surpassed 9,000 a.u. at 14 residues (+72% over 13 residues and +2,700% over the initial 6-residue case). Both the minimum and maximum energy values also expanded rapidly with sequence length. For example, minimum energy increased from approximately 10 a.u. (5 residues) to over 23,000 a.u. (14 residues), while maximum energy increased from  15 a.u. to over 32,000 a.u.—representing increases of approximately 230,000% and 210,000%, respectively. These nonlinear surges demonstrate the exponential growth of the quantum energy landscape as conformational dimensionality increases. This nonlinear energy amplification reveals two key insights. First, the quantum system is capable of encoding subtle structural variations with high precision: even the addition or removal of a single amino acid causes energy changes distinguishable by the quantum model. Second, the energy of the quantum system, as encoded under our Hamiltonian mapping, is highly correlated with the underlying conformational energy of the protein fragment. Thus, the lowest-energy quantum state produced by VQE directly corresponds to the predicted structure, validating the soundness of our prediction strategy. It is precisely due to this energy amplification behavior that we modularized the measurement stage, allowing it to operate independently with non-parametric circuits for high-precision bitstring sampling. This design preserves prediction fidelity while maintaining engineering tractability under realistic hardware constraints. It should be noted that the observed “energy amplification” trend primarily originates from the Hamiltonian construction itself, where the number and magnitude of penalty and interaction terms scale rapidly with sequence length. Quantum computation does not artificially amplify energies, but rather provides a mechanism to faithfully encode and optimize this highly nonlinear energy landscape. In this sense, the amplification reflects the exponential growth of conformational complexity under our Hamiltonian mapping, while the quantum framework enables tractable energy minimization and structural recovery under these conditions.
@@
4aoi [13] VVLPYMKHGDLRNF 14 1155–1168 102 23245.37 32378.95 9133.58 13328.65 413
4cig [52] VRDQAEHLKTAVQM 14 165–178 102 21375.59 29846.54 8470.94 17293.54 413
4tmk [39] IEGLEGAGKTTARN 14 8–21 102 22590.21 29135.42 6545.21 199292.66 413
3d7z [2] YLVTHLMGADLNNI 14 103–116 102 22979.86 29707.30 6727.43 156289.48 413
4jpx [60] DYLEAYGKGGVKA 13 154–166 92 16962.09 22231.95 5269.86 90422.62 373
1yc4 [36] ELISNSSDALDKI 13 47–59 92 16129.38 20745.81 4616.42 15777.29 373
5cqu [69] RKLGRGKYSEVFE 13 43–55 92 17865.39 22801.51 4936.12 7620.94 373
2qbs [77] HCSAGIGRSGT 11 214–224 72 6691.57 9356.87 2665.30 13899.11 293
4f5y [65] GLAWSYYIGYL 11 158–168 72 6408.50 8858.60 2450.10 6157.46 293
5cxa [61] FDGKGGILAHA 11 174–184 72 6946.43 9298.82 2352.40 5638.71 293
4mc1 [22] LLDTGADDTV 10 23–32 63 4092.24 6199.23 2106.99 5609.02 257
3d83 [2] YLVTHLMGAD 10 103–112 63 4235.34 6119.16 1883.82 19833.57 257
3vf7 [7] LLDTGADDTV 10 23–32 63 3975.02 6162.42 2187.40 5348.25 257
4y79 [6] DACQGDSGG 9 189–197 54 1549.16 2874.21 1325.05 207445.70 221
1e2l [56] AQITMGMPY 9 124–132 54 1509.66 2837.82 1328.15 12951.69 221
1gx8 [34] SAPLRVYVE 9 36–44 54 1626.02 3053.53 1427.51 14080.77 221
1m7y [4] TAGATSANE 9 117–125 54 1420.38 2714.98 1294.60 12918.04 221
3dx3 [38] HNDPGWI 7 90–96 38 339.99 962.62 622.63 4661.24 157
5c28 [67] CDLCSVT 7 663–669 38 386.81 792.78 405.97 114029.96 157
3ibi [72] IQFHFH 6 91–96 23 120.66 455.42 334.76 4486.62 97
2v25 [49] ATFTIT 6 81–86 23 100.42 340.83 240.42 22356.46 97
3ckz [12] VKDRS 5 149–153 12 10.43 14.65 4.22 5763.36 53
3eax [43] RYRDV 5 45–49 12 10.36 16.02 5.66 4028.72 53
@@
Our RMSD-based geometric comparison across 23 protein binding site fragments indicates that the quantum method achieved lower RMSD values in 18 samples, outperforming AlphaFold3 (AF3), which had better results in the remaining 5 cases (Figure S1). The average RMSD for quantum predictions was 3.33 Å, with a median of 3.53 Å, while AF3 predictions yielded an average RMSD of 3.87 Å and a median of 3.92 Å. Notably, the quantum method exhibited stronger structural fidelity in shorter sequences (5-9 residues) and in regions with higher structural flexibility. The predicted structures were used as ligands in docking experiments with a standardized receptor structure, and binding affinity were recorded over multiple runs. Results demonstrate that quantum-predicted structures achieved better binding affinities in 21 out of 23 cases (Figure S2). The average binding affinity for the quantum method was -4.38 kcal/mol, while the AF3-predicted structures averaged -4.00 kcal/mol. This suggests that quantum predictions not only exhibit superior geometric precision, but also more accurately capture the energetic characteristics of real biomolecular interactions. Furthermore, a cross-analysis of structural and functional metrics reveals that quantum-predicted structures exhibit a strong alignment between geometric and energetic consistency. Specifically, conformations with lower RMSD values tend to display more favorable binding affinities. This relationship is quantitatively supported by Pearson (0.836-0.836) and Spearman (0.823-0.823) coefficients (see Supplementary Figure S10), confirming a statistically significant correlation between RMSD and docking affinity across the tested fragments. This pattern suggests that, within the current problem scale, quantum predictions effectively capture both low-energy conformations and biologically relevant binding configurations. In summary, the quantum computing framework presented in this study delivers structural outputs on real quantum hardware that outperform state-of-the-art deep learning methods in both geometric and functional aspects. Particularly in low-information or small-sample prediction scenarios, where traditional models tend to degrade, our framework demonstrates significant advantages, highlighting its potential utility and scalability in structural biology and bioengineering applications.
In addition to the comparison with AlphaFold3, we further benchmarked our quantum framework against classical simulation–based structure prediction methods to examine its geometric and energetic advantages more broadly. As shown in Supplementary Figures S3 and S4, the quantum predictions consistently produced lower RMSD and more favorable docking affinities than their classical counterparts. Across the 23 tested fragments, the quantum RMSD averaged 3.68 Å (median 3.54 Å) versus 6.34 Å (median 6.12 Å) for classical models, reflecting a 42% improvement in geometric accuracy. Correspondingly, the mean binding affinity improved from –3.48 kcal mol-1 (classical) to –4.52 kcal mol-1 (quantum), with 21 of 23 samples exhibiting lower (better) docking free energies. The most significant geometric gains were observed in flexible or short-chain fragments such as 2qbs, 3ckz, and 4tmk, where quantum-predicted backbones maintained native-like conformations without excessive over-compaction typical of classical energy-minimization approaches. Energetically, cases like 4aoi, 4tmk, and 4y79 demonstrated binding-free-energy reductions exceeding 1 kcal mol-1, suggesting that the quantum ansatz more accurately encodes local electronic interactions relevant to ligand recognition. These trends together indicate that the proposed quantum–variational framework not only enhances structural fidelity but also captures the physical realism of binding energetics, offering a coherent description of biomolecular conformations that classical predictors often approximate empirically.
@@
Compared to AlphaFold3 predictions, the quantum method’s predicted structures achieved approximately 20% reductions in the upper and lower bounds of RMSD in different test cases during docking tests, indicating higher consistency and accuracy in structural prediction (Figure 3H-I). Furthermore, compared to the classical minimization baseline, the quantum method achieved an even larger improvement, with average RMSD upper- and lower-bound reductions of 35–45% across all cases, demonstrating that quantum optimization can escape shallow local minima that often trap classical gradient-based refinement. In terms of affinity score, the quantum method achieved an average improvement of about 10% over AlphaFold3 (Figure 3G). Relative to the classical pipeline, the quantum predictions yielded consistently stronger docking affinities (on average by 0.8–1.0 kcal/mol), indicating more favorable energetic stability of the predicted binding conformations. Additionally, the quantum method exhibited a tighter distribution of affinity scores, with significantly fewer outliers, further validating its reliability in optimizing docking performance.
Across all benchmark fragments, the quantum framework achieved the lowest RMSD values in five out of seven test cases when compared simultaneously against both AlphaFold3 and classical methods (Figure 3F). For example, in 1a9m, 3b26, and 3ans, quantum predictions achieved RMSD reductions of 25–40% relative to the classical minimization and 15–20% relative to AF3, reflecting a closer recovery of experimental structures. Only in 6mu3, due to the short peptide length and limited conformational flexibility, did all methods converge to similar results within a 0.1 Å difference.
@@
For protein structure prediction, besides evaluating predictive accuracy in docking tasks through binding affinity scores and RMSD bounds of docking poses, another important metric is the RMSD itself. The RMSD (root-mean-square deviation) metric is used to quantify structural differences between the predicted and experimental structures. In our experiments, we compared the RMSD of structures predicted by the VQE method and those predicted by AF3 against the corresponding experimental structures in the dataset. As shown in Figure 3F, the VQE method outperformed AF3 in six out of seven cases. When the classical baseline is also considered, the overall ranking of RMSD performance follows: Quantum << AF3 << Classical, confirming the progressive improvement from classical to quantum methods. For 6mu3, due to its short chain length and the limited degrees of freedom of the modeling approach, it lagged behind AF3 by only a percentage-level margin. Although most experimental results did not significantly surpass AF3 predictions due to limitations in modeling precision, they nevertheless demonstrate that quantum methods possess certain advantages.